# Characteristic Guidance: Non-linear Correction for DDPM at Large Guidance Scale

### v1

@@ 5.4 Cifar 10: Natural Image Generation
We compare the performance of classifier-free guidance and characteristic guidance on generating natural images with DDPM trained on the Cifar 10 dataset. The quality of generated image are evaluated by the Frechet Inception Distance (FID) (Heusel et al., 2017), the Inception Score (IS) (Salimans et al., 2016) and visual inspection.
Aircraft images generated by both classifier-free and characteristic guidance are compared in Fig.5 at various guidance scales. We fix the initial random noise for both classifier-free and characteristic guided DDPMs, therefore samples at small $\omega$ have same context. However, for large guidance scale $\omega$, images generated by classifier-free guidance tend to have dull or even white backgrounds, while images guided by characteristic guidance have more vivid and realistic backgrounds with sky and clouds. This means characteristic guidance is less biased in color than classifier-free guidance.
Table 1: FID and IS of characteristic guidance (CH) and classifier free guidance (CF) guided DDPM on Cifar10 dataset. Samples are drawn with DDIM and DPM++2M methods. Characteristic guidance achieves similar IS with classifier-free guidance while significantly enhance the FID when the guidance scale $\omega$ is large.
$\omega$ DDIM
FID IS
CF CH CF CH
0.3 4.52 4.46 9.18 9.23
0.6 4.80 4.64 9.39 9.43
1.0 6.22 5.86 9.52 9.53
1.5 8.56 7.89 9.50 9.57
2.0 11.02 10.10 9.49 9.51
4.0 19.85 18.15 9.23 9.22
6.0 27.04 24.77 8.85 8.86
$\omega$ DPM++2M
FID IS
CF CH CF CH
0.3 3.33 3.35 9.69 9.69
0.6 3.51 3.44 9.93 9.93
1.0 4.75 4.51 10.06 10.04
1.5 6.85 6.37 10.08 10.07
2.0 9.11 8.34 10.04 10.04
4.0 17.04 15.52 9.75 9.76
6.0 23.47 21.46 9.33 9.38
The evaluation metrics for sample quality are listed in Table.1, highlighting that characteristic guidance typically yields samples with lower FID and comparable IS to classifier-free guidance. Two fast sampling methods are tested: DDIM (50 steps), and DPM++2M (50 steps), with the operator $\mathbf{P}$ for characteristic guidance to be projection from the input to its channel-wise mean. FID assesses the discrepancy between sample and dataset distributions, whereas IS evaluate the InceptionV3 model’s classification results on these samples. Consequently, characteristic guidance appears more effective in stimulating the InceptionV3 model with features that align closely with the original dataset, causing significant FID drop.
@@ 5.5 ImageNet 256: Correction for Latent Space Model
The characteristic guidance works on latent space diffusion models (LDM) (Rombach et al., 2021) in a different way: it enhances the control strength rather than lowering the FID. We test the characteristic guidance on the LDM for ImageNet256 dataset, with codes and models adopted from (Rombach et al., 2021). The quality of generated image are evaluated by the Frechet Inception Distance (FID), the Inception Score (IS) and visual inspection.
Figure 6: Comparison of FID and IS between characteristic guidance (CH) and classifier free guidance (CF) guided DDPM on ImageNet256 dataset. Samples are drawn with DDIM and DPM++2M (not completed, it is now copied from DDIM results) methods from latent diffusion model. Characteristic guidance achieves better IS than classifier-free guidance but worse FID. However, the IS-FID curve of CH and CF guidance coincide with each other.
In Fig.1, we compare volcano images generated using both classifier-free and characteristic guidance across various guidance scales. For both methods, we maintained consistent initial random noise, ensuring that samples at smaller $\omega$ values share similar contexts. However, at larger guidance scales $\omega$, images guided by the classifier-free method show notable color distortion and underexposure. In contrast, characteristic guidance maintains color consistency and proper exposure, progressively accentuating volcanic features like smoke and lava. This comparison highlights characteristic guidance’s capability in latent space, effectively mitigating color issues, addressing exposure problems, and enhancing context control.
@@
In Fig.6, we present sample quality evaluation metrics using two fast sampling methods: DDIM (50 steps) and DPM++2M (50 steps). Characteristic guidance, while improving the Inception Score (IS), adversely impacts the Fréchet Inception Distance (FID) compared to classifier-free guidance. This IS-FID trade-off complicates the evaluation of which guidance performs better. Suprisingly, the IS-FID curves of both methods align closely, with the curve for characteristic guidance showing a rightward shift. This shift implies that characteristic guidance achieves higher IS at lower guidance scales, effectively operating at a higher effective guidance scale $\omega$ than classifier-free guidance, suggesting enhanced control strength.
@@
The differing functionality of characteristic guidance on Cifar10 and ImageNet may be attributed to our choice of projection operator $\mathbf{P}$. In latent space, we use channel-wise projection to the residual vector, not to the channel-wise mean, as the latter’s correction on RGB color does not make sense in latent space. Despite the unclear interpretation of this approach in latent space, our results suggest a significant correlation with control strength and the guidance scale $\omega$.
5.6 The Stopping Criteria for Iteration
@@
Characteristic guidance provides non-linear correction by solving (6) iteratively. Such iteration happens for every time step and is controlled by a threshold (tolerance $\eta$ of residual in Appendix.D) and a maximal iteration number (set to 10) as the stopping criteria. We exam how the threshold affects evaluation metrics (FID and IS) and the total number of iteration, which heavily affects the computational efficiency because each iteration calls the de-noising neural network.
@@
Figure 7: Convergence and iteration analysis on Cifar10 dataset ( $\omega=6$, DPM, 50 steps ). The graph shows the relationship between threshold levels and their effects on Fréchet Inception Distance (FID) and Inception Score (IS). It also plots the iteration count of characteristic guidance, for a typical generation process, across various thresholds against diffusion time step $T$, ranging from 00 (image) to $1000$ (noise).
Fig.7 illustrates the relationship between various threshold levels (parameter $\eta$ in Appendix.D) and their corresponding effects on Fréchet Inception Distance (FID) and Inception Score (IS). It is observed that convergence is typically attained at thresholds lower than $1e^{-3}$. Additionally, the chart also displays the number of iterations employed by characteristic guidance at different thresholds, plotted against the diffusion time step $T$, which varies from 00 (image) to $1000$ (noise). It shows that the non-linear corrections mainly take place at the middle of the diffusion process, which is known to be a critical stage for context generation (Zhang et al., 2023).
@@ 6 Conclusion
We introduced characteristic guidance, an innovative approach providing non-linear correction to classifier-free guided DDPMs. This method enforces adherence to the Fokker-Planck equation of the diffusion process in a theoretically robust, training-free, and derivative-free manner, while maintaining compatibility with current sampling techniques. We evaluated characteristic guidance across various scenarios, such as conditional Gaussian, mixture of Gaussian, the Landau-Ginzburg model for magnets, Cifar10, and Imagenet. The results demonstrate its effectiveness in reducing bias and enhancing control.
Furthermore, characteristic guidance has shown that it can address the shortcomings of DDPM at large guidance scales through a correction $\Delta\mathbf{x}$ as in (5). This formula could lay the theoretical groundwork for future studies aiming for large guidance scale correction, with more effective $\Delta\mathbf{x}$ beyond our characteristic guidance.
Limitations
The process of resolving the non-linear correction (6) requires iterative computations involving neural networks. This approach, without effective regularization, tends to be slow and susceptible to instability. Exploring advanced regularization techniques beyond the projection $\mathbf{P}$ could enhance the convergence rate hence accelerate the computation.
Moreover, the harmonic ansatz employed here, although less restrictive than the linear additivity assumption in classifier-free guidance, is not a good approximation for small diffusion times $T$. A more in-depth exploration of the Laplacian of score functions is necessary to refine and improve the accuracy of the correction.
@@ Appendix C Fokker-Plank Equation
Appendix D Fixed Point Iteration Methods for Solving the Correction Term
@@
0:  $\mathbf{x},\mathbf{c},t_{i},\omega,\boldsymbol{\epsilon}_{\theta},\sigma_{i}$, $\bar{\alpha}_{i}$, $\eta$(tolerance), $\gamma$(relaxation factor)
@@
13:           $\mbox{del}\mathbf{g}_{B}[0]$;  $\mbox{del}\Delta\mathbf{x}_{B}[0]$;
@@
Appendix E Deriving the Characteristic Guidance Using the Method of Characteristics
It is not easy to express $\mathbf{s}(\mathbf{x},t)$ as a function of $\mathbf{s}_{1}(\mathbf{x},t)$ and $\mathbf{s}_{2}(\mathbf{x},t)$ without training or taking complex derivatives. Fortunately, the harmonic ansatz 4.1 leads us to a viable solution. The harmonic ansatz reduces the score Fokker-Planck equation (31) to the following non-linear first-order partial differential equation (PDE)
Characteristic Lines
where $\mathbf{x}_{0}=\mathbf{x}(0)$ is the starting point of the line and serves as a constant parameter, $\mathbf{s}(\mathbf{x}_{0},0)$ depends on the initial condition of $\mathbf{s}$ at $t=0$. An important remark on the characteristic lines is that (37) is valid for $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ as well, if we replace every $\mathbf{s}$ with $\mathbf{s}_{1}$ or $\mathbf{s}_{2}$. This is because $\mathbf{s}$, $\mathbf{s}_{1}$, and $\mathbf{s}_{2}$ are solutions to the same PDE (34). The merit of these characteristic lines lies in their ability to determine the value of $\mathbf{s}(\mathbf{x}(t),t)$ at $t>0$ using information of $\mathbf{x}_{0}$ and $\mathbf{s}(\mathbf{x}_{0},0)$ at $t=0$, thereby allowing us to utilize the initial condition (33).
Deducing $\mathbf{s}(\mathbf{x},t)$
$\mathbf{x}(0)=\mathbf{x}_{1}(0)=\mathbf{x}_{2}(0)=\mathbf{x}_{0},$ (38)
then the initial condition (33) and the equation (37) tells that
Simplify the System
$(1+\omega)\mathbf{x}_{1}-\omega\mathbf{x}_{2}=\mathbf{x}.$ (42)
This inspires us to define the correction term $\Delta\mathbf{x}$ as
Characteristic Guidance and Lemma 4.2
and
@@ Appendix F Conditional Gaussian
The score function of the unconditional, conditional, and guided model along the forward diffusion process can be solved analytically from the Fokker-Planck equation:
@@ Appendix H Diffusion model for Science: Cooling of Magnet
Guidance, as described in 27, can be conceptualized as a cooling process with the guidance scale $\omega$ representing the temperature drop. We demonstrate that the guidance can simulate the cooling process of a magnet around the Curie temperature, at which a paramagnetic material gains permanent magnetism through phase transition. We also demonstrate that non-linear correction for classifier-free guidance is important for correctly characterize the megnet’s phase transition.
@@
Appendix I Cifar 10
Appendix J Latent Diffusion Model

### v5

@@ 6.3 CIFAR-10: Natural Image Generation
In this study, we evaluate the effectiveness of classifier-free (CF) and characteristic (CH) guidance in generating natural images using a DDPM trained on the CIFAR-10 dataset. We assess the quality of the generated images using Frechet Inception Distance (FID), Inception Score (IS), and visual inspection (Sec 6.6).
Table 1: Comparison of FID and IS for CF and CH guided DDPM on the CIFAR-10 dataset, using DDIM (50 steps) and DPM++2M (50 steps). The operator $\mathbf{P}$ for CH is the projection from the input to its channel-wise mean. Characteristic guidance generally achieves lower FID and comparable IS, particularly at higher guidance scales $\omega$.
$\omega$ DDIM DPM++2M
FID $\downarrow$ IS $\uparrow$ FID $\downarrow$ IS $\uparrow$
CF CH CF CH CF CH CF CH
0.3 4.52 4.46 9.18 9.23 3.33 3.35 9.69 9.69
0.6 4.80 4.64 9.39 9.43 3.51 3.44 9.93 9.93
1.0 6.22 5.86 9.52 9.53 4.75 4.51 10.06 10.04
1.5 8.56 7.89 9.50 9.57 6.85 6.37 10.08 10.07
2.0 11.02 10.10 9.49 9.51 9.11 8.34 10.04 10.04
4.0 19.85 18.15 9.23 9.22 17.04 15.52 9.75 9.76
6.0 27.04 24.77 8.85 8.86 23.47 21.46 9.33 9.38
Our experiments, detailed in Table.1, demonstrate that characteristic guidance achieves a lower FID while maintaining a comparable IS to classifier-free guidance, illustrating its effectiveness in balancing control strength and diversity. Unlike the typical trade-off between FID and IS in guidance methods (Ho, 2022), where enhancing diversity marked by improved FID often reduces control strength by a decrease in IS, characteristic guidance improves FID without compromising control strength. These results underscore its advantage in producing high-fidelity images while preserving both diversity and control.
@@ 6.4 ImageNet 256: Correction for Latent Space Model
The characteristic guidance’s effect on latent space diffusion models (LDM) (Rombach et al., 2021) is different: it enhances the control strength rather than lowering the FID. We test the characteristic guidance on the LDM for ImageNet256 dataset, with codes and models adopted from the LDM paper (Rombach et al., 2021). The quality of generated images are evaluated by FID, IS and visual inspection (Sec 6.6).
Figure 6: FID and IS comparison between CH and CF guided DDPM on ImageNet 256, using DDIM and DPM++2M sampling methods. CH shows better IS but higher FID than CF, with a rightward shift in its IS-FID curve indicating higher effective guidance scale.
@@
Fig.6 presents sample quality metrics using DDIM (50 steps) and DPM++2M (50 steps). While CH improves IS, it adversely affects FID compared to CF. This IS-FID trade-off complicates performance evaluation. Notably, the IS-FID curves of CH and CF align closely, with CH demonstrating a rightward shift. This means CH guidance attains a certain IS with lower guidance scale, suggesting enhanced control strength.
@@
The difference in CH guidance’s functionality between CIFAR-10 and ImageNet256 may relate to the choice of projection operator $\mathbf{P}$. In latent space, we apply channel-wise projection to the residual vector, avoiding channel-wise mean projection inappropriate for RGB color correction in latent space. Despite the unclear interpretation of this approach in latent space, our visual inspection suggest CH guidance achieves semantic characteristics enhancing and exposure correction, as shown in Fig.5 (Right).
6.5 Stable Diffusion: Text/Image to Image Generation
Characteristic guidance, designed for broad compatibility, can be seamlessly integrated with any DDPM that supports classifier-free guidance. We have successfully integrated our characteristic guidance into Stable Diffusion WebUI (AUTOMATIC1111, 2023) as public available extension, supporting all provided samplers under both Txt2Img and Img2Img mode. Fig.1 showcases a comparative visualization of images from Stable Diffusion XL (Podell et al., 2023). More visulizations could be found in Appendix.
6.6 Visual Inspection
This section evaluates images generated using classifier-free (CF) and characteristic (CH) guidance from the same initial noise, demonstrating that CH guidance is more capable in enhancing semantic characteristics.
For aircraft images created by the Cifar-10 model, as shown in Fig.5 (Left), CF guidance at higher scales ($\omega$) produces backgrounds that are often dull or white. In contrast, CH guidance results in more vivid and realistic backgrounds, such as skies with clouds that open appears in aircrafts photos. Similarly, with volcano images from ImageNet 256’s latent diffusion models (Fig.5, Right), CF-guided images at elevated $\omega$ levels exhibit color casting and underexposure. CH guidance, conversely, maintains consistent color and exposure, better accentuating key volcanic elements like smoke and lava, thus emphasizing the semantic characteristics more effectively.
In stable diffusion scenarios, Fig.1 compares CF and CH guidance across various prompts and CFG scales. At a CFG scale of 10 ($\omega=9$), CH guidance mitigates anatomical irregularities, as seen in the handstand girl images (Fig.1 Left). Increasing the scale to 30 ($\omega=29$), CF guidance struggles with color cast and exposure issues, while CH guidance produces images that more aptly emphasize prompt-relevant features, such as running girls and lush mountains (Fig.1 Right).
Overall, these comparisons underscore CH guidance’s capability in not just adjusting basic elements like color and exposure, but more importantly, in amplifying the semantic characteristics of the subjects in line with the given prompts.
7 Discussion
7.1 The Stopping Criteria for Iteration.
@@ 7 Discussion
Characteristic guidance applies non-linear correction at each time step by iteratively solving (11). The iteration process is governed by two key stopping criteria: a predefined threshold (tolerance $\eta$, detailed in Appendix.E) and a maximum iteration limit set at 10. We investigate the impact of the threshold on evaluation metrics (FID and IS) and the total number of iterations, which is crucial for computational efficiency as each iteration involves calling the de-noising neural network.
As illustrated in Fig.7, we observe that convergence typically occurs at thresholds below $1e^{-3}$. This finding informs the optimal setting of $\eta$ to balance quality and efficiency. Additionally, the graph also highlights the iteration frequency employed by characteristic guidance across two threshold levels. It is noteworthy that non-linear corrections predominantly occur in the middle stages of the diffusion process, a phase critical for context generation as identified in (Zhang et al., 2023).
@@
Figure 7: Analysis of convergence and iterations in characteristic guidance on the CIFAR-10 dataset ($\omega=6$, DPM++2M, 50 steps). The graph depicts the influence of varying threshold levels on FID and IS, and the iteration count across different thresholds relative to the diffusion time step $T$, ranging from 00 (image) to $1000$ (noise).
7.2 The Difference with Dynamical Thresholding
Dynamical thresholding (Saharia et al., 2022) is a technique designed for RGB images. It predicts the final output image $\mathbf{x}_{0}$ with (1) at each step, correcting and normalizing any out-of-range pixels back to $[-1,1]$.
Dynamical threshold is effective in correct color cast when pixel values representing RGB colors. But it falls short in general tasks like latent space sampling and correcting contextual inaccuracies. In contrast, characteristic guidance is a theoretical method that is applicable to any continuous data type of any dimension. In practice, as Fig.20 shows, characteristic guidance is capable of correcting contextual inaccuracies while dynamical thresholding focuses more on color.
@@ 8 Conclusion
We introduced characteristic guidance, a novel guidance method providing non-linear correction to classifier-free guided DDPMs. It is, as far as we know, the first attempt to lay the theoretical framework for large guidance scale correction. Our comprehensive evaluations of characteristic guidance covered various scenarios, each with a distinct objective: validating its theoretical advantage on Gaussian models, applying it to physics problems like the Landau-Ginzburg model for magnetism, and demonstrating its robustness in image generation across different spaces — CIFAR-10 in pixel-space, ImageNet and Stable Diffusion in latent space. The results consistently highlight the method’s effectiveness in enhancing the semantic characteristics of subjects and mitigating generation irregularities.
9 Limitation
FID’s Effectiveness. The effectiveness of FID for evaluating guided DDPM is compromised by the difference between the target conditional probabilities $p(\mathbf{x}|\mathbf{c},\omega)$ and marginal probabilities $p(\mathbf{x})$. FID’s focus on the distance from $p(\mathbf{x})$ contrasts with guided DDPM’s goal of sampling from $p(\mathbf{x}|\mathbf{c},\omega)$, suggesting the need for more apt metrics that address this discrepancy.
Speed. The process of resolving the non-linear correction (11) requires iterative computations involving neural networks. This approach, without effective regularization, tends to be slow. Exploring advanced regularization techniques beyond the projection $\mathbf{P}$ could enhance the convergence rate hence accelerate the computation.
Software and Data
Characteristic guidance is available at https://scraed.github.io/CharacteristicGuidance/ as open source and public extension for Stable Diffusion WebUI (AUTOMATIC1111, 2023), supporting all provided samplers under both Txt2Img and Img2Img mode.
Acknowledgements
We would like to express our gratitude to Professor Yang Wang at the Hong Kong University of Science and Technology for his support of this study.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix C Fokker-Plank Equation
Note that when the diffusion time steps are infinitesimally small, equation (43) could be rewritten in terms of $\boldsymbol{\epsilon}$ with the help of the relation (33).
$\frac{\partial\boldsymbol{\epsilon}}{\partial t}=\frac{1}{2}\left(\nabla_{\mathbf{x}}(\boldsymbol{\epsilon}\cdot\mathbf{x})+\nabla_{\mathbf{x}}^{2}\boldsymbol{\epsilon}-\frac{\nabla_{\mathbf{x}}\|\boldsymbol{\epsilon}\|^{2}_{2}}{\sqrt{1-e^{-t}}}+\frac{e^{-t}}{(1-e^{-t})}\boldsymbol{\epsilon}\right)$ (45)
@@
Appendix D Deriving the Characteristic Guidance Using the Method of Characteristics
It is not easy to express $\mathbf{s}(\mathbf{x},t)$ as a function of $\mathbf{s}_{1}(\mathbf{x},t)$ and $\mathbf{s}_{2}(\mathbf{x},t)$ without training or taking complex derivatives in Eq. (4). Fortunately, the harmonic ansatz 5.1 leads us to a viable solution. The harmonic ansatz reduces the score Fokker-Planck equation (4) to the following non-linear first-order partial differential equation (PDE)
Characteristic Lines
where $\mathbf{x}_{0}=\mathbf{x}(0)$ is the starting point of the line and serves as a constant parameter, $\mathbf{s}(\mathbf{x}_{0},0)$ depends on the initial condition of $\mathbf{s}$ at $t=0$. Besides, such characteristic lines of $\mathbf{s}$ also induces characteristic lines of $\boldsymbol{\epsilon}$ as they are connected by (33).
An important remark on the characteristic lines is that (50) is valid for $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ as well, if we replace every $\mathbf{s}$ with $\mathbf{s}_{1}$ or $\mathbf{s}_{2}$. This is because $\mathbf{s}$, $\mathbf{s}_{1}$, and $\mathbf{s}_{2}$ are solutions to the same PDE (47). The merit of these characteristic lines lies in their ability to determine the value of $\mathbf{s}(\mathbf{x}(t),t)$ at $t>0$ using information of $\mathbf{x}_{0}$ and $\mathbf{s}(\mathbf{x}_{0},0)$ at $t=0$, thereby allowing us to utilize the initial condition (46).
Deducing $\mathbf{s}(\mathbf{x},t)$
$\mathbf{x}(0)=\mathbf{x}_{1}(0)=\mathbf{x}_{2}(0)=\mathbf{x}_{0},$ (51)
then the initial condition (46) and the equation (50) tells that
Simplify the System
$(1+\omega)\mathbf{x}_{1}-\omega\mathbf{x}_{2}=\mathbf{x}.$ (55)
This inspires us to define the correction term $\Delta\mathbf{x}$ as
Characteristic Guidance, Lemma 5.2, and Theorem 5.3
and
As for Theorem 5.3, it is a direct consequence of Lemma 5.2. Lemma 5.2 says that $\mathbf{s}(\mathbf{x},t)$ is a solution of the equation (47), then its corresponding $\boldsymbol{\epsilon}(\mathbf{x},t)$, automatically satisfies the FP equation (4) except the Laplacian term, which can be further eliminated by the Harmonic ansatz. Therefore $\boldsymbol{\epsilon}(\mathbf{x},t)$ in (59) has no mixing error (7) under the Harmonic ansatz.
Appendix E Fixed Point Iteration Methods for Solving the Correction Term
@@
0:  $\mathbf{x},\mathbf{c},t_{i},\omega,\boldsymbol{\epsilon}_{\theta},\sigma_{i}$, $\eta$(tolerance)
@@
13:           $\mbox{delete}\ \mathbf{g}_{B}[0]$;  $\mbox{delete}\ \Delta\mathbf{x}_{B}[0]$;
@@ Appendix F Conditional Gaussian
The score function of the unconditional, conditional, and guided model along the forward diffusion process can be solved analytically from the FP equation:
@@ Appendix H Diffusion Model for Science: Cooling of Magnet
Guidance, as described in 39, can be conceptualized as a cooling process with the guidance scale $\omega$ representing the temperature drop. We demonstrate that the guidance can simulate the cooling process of a magnet around the Curie temperature, at which a paramagnetic material gains permanent magnetism through phase transition. We also demonstrate that non-linear correction for classifier-free guidance is important for correctly characterize the megnet’s phase transition.
@@
Appendix I Experimental Details on Cifar-10 and ImageNet 256
Our CIFAR-10 experiment utilized a modified U-Net model with 39 million parameters, sourced from a classifier-free diffusion guidance implementation (Coderpiaobozhe, 2023). The model featured a channel size of 128 and a conditional embedding dimension of 10, operating over 1000 timesteps. Training involved 400,000 iterations with a batch size of 256, using the AdamW optimizer at a $2\times 10^{-4}$ learning rate and a 0.1 dropout rate. Stability was ensured with an EMA decay rate of 0.999, and the diffusion process followed a linear $\beta_{t}$ schedule from $10^{-4}$ to $0.02$ over 1000 timesteps. The characteristic guidance is computed by Alg.2 with $\gamma=0.002$, $\eta=0.001$, and $\alpha=0.999$. All CIFAR-10 experiments were conducted on 1 NVIDIA Geforce RTX 3090 GPU.
For the ImageNet 256 experiment, we utilized a pretrained latent diffusion model, as detailed in (Rombach et al., 2021), featuring over 400 million parameters. This model is configured to run with a total of 1000 timesteps and works on latents with the shape of $3\times 64\times 64$. The characteristic guidance is computed by Alg.3 with $\gamma=1$, $m=2$, $\eta=0.005$. Experiments were conducted on 6 NVIDIA Geforce RTX 4090 GPUs.
Appendix J Visualization Results on Latent Diffusion Model